INSTRUCTIONS
A. Prepare a tabular summary of transactions
B. Formulate the income statement, retained earnings statement and balance sheet as

of September 31st, 2018

Information:

After 2 years of managing a established business, INNO reports an annual growth margin of 38%; for the term ended in September 31st, 2018, they reported the following transactions:

  1. Stockholders invested $10,000 in cash in exchange for common stock of INNOmuebles

  2. Paid $800 for July rent on office space.

  3. Purchased office equipment on account $3,000.

  4. Sold various merchandise (furniture) to clients for cash $1,500.

  5. Borrowed $700 cash from a bank on a note payable.

  6. Sold office furniture to a specific client on account $2,000.

  7. Paid monthly expenses: salaries $500, utilities $300, and advertising $100.
Answer #1

A. Prepare a tabular summary of transactions

Table for Summary of transactions
S.No Transactions Assets = Liability + Capital
Cash/Bank + Stock + Debtors Total Loan + Accounts payable Total
1 Investment 10000 + 0 + 0 10000 = 0 + 0 0 10000
2 Purchase of Stock 0 + 830 + 0 830 = 0 + 3000 3000 -2170
3 Sales on cash 1500 + 0 + 0 1500 = 0 + 0 0 1500
4 Rent -800 + 0 + 0 -800 = 0 + 0 0 -800
5 Borrowing 700 + 0 + 0 700 = 700 + 0 700 0
6 Sales on account 0 + 0 + 2000 2000 = 0 + 0 0 2000
7 Salary -500 + 0 + 0 -500 = 0 + 0 0 -500
8 Utilites -300 + 0 + 0 -300 = 0 + 0 0 -300
9 Advertisement -100 + 0 + 0 -100 = 0 + 0 0 -100
10500 + 830 + 2000 13330 = 700 + 3000 3700 9630

B. Formulate the income statement, retained earnings statement and balance sheet as

of September 31st, 2018

Income statement as at September 30, 2018
Particulars Amount Particulars Amount
To Purchase 3000 By Sales
-Cash 1500
To Gross Margin 1330 -on account 2000
By Closing stock 830
4330 4330
To Salary 500 By Gross Margin 1330
To Utilites 300
To Advertisement 100
Net profit 430
Retained earning as at September 30, 2018
Particulars Amount
Opening Balance Nil
Profit during the period 430
Closing Balance 430
Balance sheet as on September 30, 2018
Liability/Capital Amount Assets Amount
Capital 9630 Cash 10500
Accounts payable 3000 Stock 830
Loan 700 Debtors 2000
Answer #2
Add you answer here...